OpenStudy (anonymous):

What number is 315% of 26.2?

OpenStudy (anonymous):

What is 315 percent of 26.2 - step by step solution Simple and best practice solution for 315% of 26.2. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the calculator fields your own values, and You will get the solution. What is % of ? is what percent of ? To get the solution, we are looking for, we need to point out what we know. 1. We assume, that the number 26.2 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 26.2 is 100%, so we can write it down as 26.2=100%. 4. We know, that x is 315% of the output value, so we can write it down as x=315%. 5. Now we have two simple equations: 1) 26.2=100% 2) x=315% where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 26.2/x=100%/315% 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for what is 315% of 26.2 26.2/x=100/315 (26.2/x)*x=(100/315)*x - we multiply both sides of the equation by x 26.2=0.31746031746*x - we divide both sides of the equation by (0.31746031746) to get x 26.2/0.31746031746=x 82.53=x x=82.53 now we have: 315% of 26.2=82.53

OpenStudy (texaschic101):

remember the percent formula that I showed you... percent formula : is/of = %/100 what number is 315% of 26.2 " is " = x " of " = 26.2 % = 315 now sub all your stuff in x/26.2 = 315/100 -- cross multiply 100x = (26.2)(315) 100x = 8253 x = 8253/100 x = 82.53 check.. what number is 315% of 26.2...turn percent to decimal 82.53 = 3.15(26.2) 82.53 = 82.53 (correct) your number is : 82.53